Let "f(x)=e^{-7x^2}" .

Then f(x) has a relative minimum at

x=


a relative maximum at

x=


and inflection points at

x=


and at

x=


Write DNE if any of the above do not exist. Write the inflection points (if any) in numerical order, smallest first.


1
Expert's answer
2020-12-10T20:24:24-0500

Solution:Domain of given function is "-\infty \:<x<\infty"

"\mathrm{Combine\:the\:critical\:point\left(s\right):}\:x=0\:\mathrm{with\:the\:domain}"

"f(x)=e^{-7x^2}"

"f'(x)=-14xe^{-7x^2}"

"-14xe^{-7x^2}=0"

"x=0"


"\mathrm{The\:function\:monotone\:intervals\:are:}" "-\infty \:<x<0" , "0<x<\infty \:"

We should check the sign "f'(x)=-14xe^{-7x^2}" at each monotone interval

"Summary\:of\:the\:monotone\:intervals\:behavior" :


1) "-\infty \:<x<0\Rightarrow increasing\Rightarrow+"

2) "0<x<\infty \:" "\Rightarrow decreasing\Rightarrow -"

3) "x=0 \Rightarrow max \Rightarrow0"

"\mathrm{Plug\:the\:extreme\:point}\:x=0\:\mathrm{into}\:e^{-7x^2}\quad \Rightarrow \quad \:y=1"


f(x) has a relative minimum at:DNErelative maximum at:

"\mathrm{maximum}\left(0,\:1\right)"


inflection points at:

"\mathrm{If\:}f\:''\left(x\right)>0\mathrm{\:then\:}f\left(x\right)\mathrm{\:concave\:upwards.}"

"\mathrm{If\:}f\:''\left(x\right)<0\mathrm{\:then\:}f\left(x\right)\mathrm{\:concave\:downwards.}"


"f''(x)=0"


x="\left(-\frac{\sqrt{14}}{14},\:\frac{1}{\sqrt{e}}\right)"


and at

x="\left(\frac{\sqrt{14}}{14},\:\frac{1}{\sqrt{e}}\right)"